# New Projection Operators for Planar Electrodynamics

### v1

@@ 1  Introduction
In Quantum Field Theory (QFT) the calculation of the propagator is essential for the perturbative development in obtaining cross sections [zuber, schwartz, peskin], as well as for elucidating the nature of the excitations that arise in the interaction between particles and fields [accioly, baeta, boldo]. For this reason, it becomes important to know different methods for calculating these propagators.
It is well known that the pole structure of the propagator encodes the physical dispersion relations of the excitations, since the on-shell energies follow from the zeros of the corresponding denominators in momentum space. Likewise, the analysis of the residues at these poles-in particular, the sign of the residue in the saturated amplitude-provides a direct criterion for tree-level unitarity and is closely tied to the causal properties of the propagating modes [accioly, baeta, boldo].
In this work we present an alternative method to obtain the propagator of three-dimensional models, in particular the case of Maxwell-Lee-Wick-Chern-Simons [leewick] and Maxwell-Deser-Jackiw electrodynamics [djackiw]. The proposed method consist of the decomposition into direct sums of subspaces; this method introduce a new set of projection operators that make easier the analysis of causality and unitarity of the propagator.
We begin by presenting a general three-dimensional theory for a vectorial field, the problem in Section 2, exemplified by the case of Maxwell-Lee-Wick-Chern-Simons electrodynamics [leewick], taking as a starting point the Lagrangian and obtaining the corresponding field equations in terms of the wave operator. In the Section 3, we review some concepts and theorems of Linear Algebra, whose demonstrations are found in the Appendix. This allows us to introduce the method and present its implications in a clearer way, in addition to its application. Next, in the Section 4, we apply the method to Electrodynamics. Later, in the Section 5, we apply the method to obtain the propagator of Maxwell-Lee-Wick-Chern-Simons and Maxwell-Deser-Jackiw electrodynamics. In the Section 6, we present the causality and unitarity analysis for the models. Finally, in the Section 7, we summarize our main results and discuss possible extensions and future directions of the proposed work.
@@ 3  Mathematical Formalism
The discussion below is grounded on a general algebraic result, namely the Cayley-Hamilton theorem [lima_algebra, axler2024]. Since the vector space $V$ is finite-dimensional, any linear operator acting on it satisfies its own characteristic polynomial. When such an operator is invertible, this implies that its inverse can be written as a linear combination of the identity and its powers. In the context considered here, this theorem provides the conceptual justification for the closure properties assumed for the operators involved and explains why propagators can be consistently expressed within the same algebraic framework.
@@ 4  Solutions of the Problem
Since the wave operator in momentum space (1), where $\alpha(p),\beta(p),\gamma(p)\in\mathbb{C}$. Our main goal will be to invert it in the most general way possible. For this purpose, we shall begin a deeper study of the operators $\omega,\theta,S:\mathbb{C}^{3}\to\mathbb{C}^{3}$, defined by $(\omega x)^{\mu}=\omega^{\mu}_{\;\nu}x^{\nu}$, $(\theta x)^{\mu}=\theta^{\mu}_{\;\nu}x^{\nu}$ and $(Sx)^{\mu}=S^{\mu}_{\;\nu}x^{\nu}$, where $x=(x^{0},x^{1},x^{2})\in\mathbb{C}^{3}$, in order to obtain properties that will facilitate the obtaining of $\mathcal{O}^{-1}$.
@@ 4.1  Operators $\omega$, $\theta$ and $S$
Furthermore, it should be noted that $\omega$ and $\theta=I-\omega$ are orthogonal projectors that satisfy the following relations:
@@
From theorems 3.1 and 3.2 we have $\mathrm{Im}\,\omega=\ker\theta$, $\ker\omega=\mathrm{Im}\,\theta$ and the decomposition $\mathbb{C}^{3}=\mathrm{Im}\,\omega+\mathrm{Im}\,\theta$.
@@ 5  Calculation of the Propagators
With such operators, one can construct the following multiplication table.
Table 2: Products between $P_{1}$, $P_{2}$ and $P_{3}$.
$P_{1}$ $P_{2}$ $P_{3}$
$P_{1}$ $P_{1}$ 0 0
$P_{2}$ 0 $P_{2}$ 0
$P_{3}$ 0 0 $P_{3}$
@@ 5.2  The Maxwell-Lee-Wick-Chern-Simons Propagator
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4m^{2}}F_{\mu\nu}\Box F^{\mu\nu}$
@@
where $A^{\mu}$ is the vector potential, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field tensor and $J_{\mu}$ an external current. The gauge-fixing term, with the parameter $\xi$, must be included in order to make the calculation of the propagator possible. In fact, as will be shown, without this term the wave operator would not be invertible. In this way, using the coefficients of corresponding wave operator,
$\alpha=-\frac{p^{2}}{\xi},\quad\beta=\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)\quad\text{and}\quad\gamma=i\kappa,$ (17)
@@
$\mathcal{O}^{-1}=-\frac{\xi}{p^{2}}P_{1}+\frac{1}{\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)+\kappa\sqrt{p^{2}}}P_{2}$
$\frac{1}{\left(-p^{2}+\frac{p^{4}}{m^{2}}\right)-\kappa\sqrt{p^{2}}}P_{3}.$ (18)
Thus, at the base $\{\omega,\theta,S\}$, we have
$\mathcal{O}^{-1}_{\mu\nu}=-\frac{p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)}{\,p^{2}\!\left[p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}\right]}\,\theta_{\mu\nu}$
$\frac{i\,\kappa}{\,p^{2}\!\left[p^{2}\!\left(1-\frac{p^{2}}{m^{2}}\right)^{2}-\kappa^{2}\right]}\,S_{\mu\nu}-\frac{\xi}{p^{2}}\omega_{\mu\nu}\,.$ (19)
@@ 5.3  The Maxwell-Deser-Jackiw Propagator
Another interesting model can be built up from the Maxwell system by adding the higher derivative Chern-Simons extension proposed by [deserjackiw]
@@
At the base $\{\omega,\theta,S\}$, with $\gamma=ip^{2}/m$, we have
@@ 6  General analysis of causality and unitarity
6.1   Analysis on the base $\{P_{1},P_{2},P_{3}\}$
Let $J_{\mu}(p)$ be an external conserved current, the saturated amplitude is then
$A(p)=J^{*\mu}(p)\,\mathcal{O}^{-1}_{\mu\nu}(p)\,J^{\nu}(p).$ (24)
Using the form (14), we obtain
$A(p)=\frac{\|J^{(+)}\|^{2}}{\beta(p)-i\gamma(p)\sqrt{p^{2}}}+\frac{\|J^{(-)}\|^{2}}{\beta(p)+i\gamma(p)\sqrt{p^{2}}},$ (25)
where the norms $\|J^{(\pm)}\|^{2}=(J^{(\pm)})^{*}\!\cdot J^{(\pm)}\geq 0$ are taken in the physical transverse subspace. Thus, we will call the coefficient of each term in (25) as follows
@@ 6.1.1  Pole structure and general spectral analysis
Therefore all physical transverse poles are located at momentum $p^{\mu}=p_{*}^{\mu}$ satisfying
@@
For each fixed spatial momentum $\vec{p}$ these correspond to energies $p^{0}=E_{n}(\vec{p})$ given by the solutions of
@@
If $E_{n}(\vec{p})$ is real, we obtain a propagating mode with dispersion relation determined implicitly by $D(E_{n},\vec{p})=0$. For such modes, causality at the microscopic level requires that the group velocity
$v_{g}=\left|\frac{\partial E_{n}}{\partial\vec{p}}\right|\leq 1,$ (32)
@@
$\frac{\partial D}{\partial p^{0}}\Big|_{p_{*}}\,dE_{n}+\frac{\partial D}{\partial\vec{p}}\Big|_{p_{*}}\!\cdot d\vec{p}=0\quad\Rightarrow\quad\frac{\partial E_{n}}{\partial\vec{p}}=-\,\frac{\partial D/\partial\vec{p}}{\partial D/\partial p^{0}}\Bigg|_{p_{*}}.$ (33)
@@
If some solutions $E_{n}(\vec{p})$ are complex, they occur in conjugate pairs by hermiticity of the underlying Lagrangian. Writing $E_{n}(\vec{p})=\Omega_{n}(\vec{p})-i\Gamma_{n}(\vec{p})$ with $\Gamma_{n}(\vec{p})>0$, the corresponding Fourier modes behave as $\exp[-i\Omega_{n}t]\,\exp[-\Gamma_{n}t]$ and can be interpreted this later on.
@@ 6.1.2  Residues and tree–level unitarity
$D_{\pm}(p^{0},\vec{p})\simeq\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}}\,\frac{1}{p^{0}-E_{*}}.$ (36)
Therefore the residues of the saturated amplitudes are
$\text{Res}\,A_{\pm}\Big|_{p^{0}=E_{*}}=\|J^{(\pm)}\|^{2}\,\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}},$ (37)
with $\|J^{(\pm)}\|^{2}\geq 0$. For real poles and real coefficients $\alpha(p)$, $\beta(p)$, $\gamma(p)$ on the real axis, the two residues are complex conjugates,
$\bigl(\text{Res}\,A_{+}\bigr)^{*}=\text{Res}\,A_{-}.$ (38)
Tree–level unitarity requires that the terms of (37) be non-negative.
6.1.3  Equivalent analysis in the base $\{\omega,\theta,S\}$
@@
$A(p)=J^{*}\!\cdot\mathcal{O}^{-1}\!\cdot J=\frac{\beta(p)}{D(p)}\,J^{*}\!\cdot J-\frac{\gamma(p)}{D(p)}\,J^{*}\!\cdot S\!\cdot J.$ (39)
@@
$J^{*}\!\cdot S\!\cdot J=(J^{(+)})^{*}\!\cdot SJ^{(+)}+(J^{(-)})^{*}\!\cdot SJ^{(-)}=$
$i\sqrt{p^{2}}\,\|J^{(+)}\|^{2}+i\sqrt{p^{2}}\,\|J^{(-)}\|^{2}.$ (42)
@@
$A(p)=\frac{1}{D(p)}\Bigl[\bigl(\beta(p)+i\gamma(p)\sqrt{p^{2}}\bigr)\,\|J^{(+)}\|^{2}\Bigr]$
@@
we recover exactly  (25). Thus the decomposition and the analysis of poles, residues, unitarity and causality are entirely equivalent in the $\{P_{1},P_{2},P_{3}\}$ and $\{\omega,\theta,S\}$ bases. The projector formalism therefore provides a compact and geometrically transparent framework for studying higher–derivative planar gauge models at the level of their propagators.

### v2

@@ 1  Introduction
In Quantum Field Theory (QFT) the calculation of the propagator is essential for the perturbative development in obtaining cross sections [6, 11, 9], as well as for elucidating the nature of the excitations that arise in the interaction between particles and fields [1, 10, 3]. For this reason, it becomes important to know different methods for calculating these propagators.
It is well known that the pole structure of the propagator encodes the physical dispersion relations of the excitations, since the on-shell energies follow from the zeros of the corresponding denominators in momentum space. Likewise, the analysis of the residues at these poles-in particular, the sign of the residue in the saturated amplitude-provides a direct criterion for tree-level unitarity and is closely tied to the causal properties of the propagating modes [1, 10, 3].
In this work we present an alternative method to obtain the propagator of three-dimensional models, in particular the case of Maxwell-Lee-Wick-Chern-Simons [7] and Maxwell-Deser-Jackiw electrodynamics [4]. The proposed method consist of the decomposition into direct sums of subspaces; this method introduce a new set of projection operators that make easier the analysis of causality and unitarity of the propagator.
We begin by presenting a general three-dimensional theory for a vectorial field, the problem in Section 2, exemplified by the case of Maxwell-Lee-Wick-Chern-Simons electrodynamics [7], taking as a starting point the Lagrangian and obtaining the corresponding field equations in terms of the wave operator. In the Section 3, we review some concepts and theorems of Linear Algebra, whose demonstrations are found in the Appendix. This allows us to introduce the method and present its implications in a clearer way, in addition to its application. Next, in the Section 4, we apply the method to Electrodynamics. Later, in the Section 5, we apply the method to obtain the propagator of Maxwell-Lee-Wick-Chern-Simons and Maxwell-Deser-Jackiw electrodynamics. In the Section 6, we present the causality and unitarity analysis for the models. Finally, in the Section 7, we summarize our main results and discuss possible extensions and future directions of the proposed work.
@@ 3  Mathematical Formalism
The discussion below is grounded on a general algebraic result, namely the Cayley-Hamilton theorem [8, 2]. Since the vector space $V$ is finite-dimensional, any linear operator acting on it satisfies its own characteristic polynomial. When such an operator is invertible, this implies that its inverse can be written as a linear combination of the identity and its powers. In the context considered here, this theorem provides the conceptual justification for the closure properties assumed for the operators involved and explains why propagators can be consistently expressed within the same algebraic framework.
@@ 4  Solutions of the Problem
Since the wave operator in momentum space (1), where $\alpha(p),\beta(p),\gamma(p)\in\mathbb{C}$. Our main goal will be to invert it in the most general possible way. For this purpose, we shall begin a deeper study of the operators $\omega,\theta,S:\mathbb{C}^{3}\to\mathbb{C}^{3}$, defined by $(\omega x)^{\mu}=\omega^{\mu}_{\;\nu}x^{\nu}$, $(\theta x)^{\mu}=\theta^{\mu}_{\;\nu}x^{\nu}$ and $(Sx)^{\mu}=S^{\mu}_{\;\nu}x^{\nu}$, where $x=(x^{0},x^{1},x^{2})\in\mathbb{C}^{3}$, in order to obtain properties that will facilitate the obtaining of $\mathcal{O}^{-1}$.
@@ 4.1  Operators $\omega$, $\theta$ and $S$
Furthermore, it should be noted that $\omega$ and $\theta=I-\omega$ are orthogonal projectors. Thus, from theorems 3.1 and 3.2 we have $\mathrm{Im}\,\omega=\ker\theta$, $\ker\omega=\mathrm{Im}\,\theta$, $\mathbb{C}^{3}=\mathrm{Im}\,\omega\oplus\mathrm{Im}\,\theta$ and the following relations are satisfied
@@ 5.2  The Maxwell-Lee-Wick-Chern-Simons Propagator
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{\mu}{4}F_{\mu\nu}\Box F^{\mu\nu}$
@@
where $A^{\mu}$ is the vector potential, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field tensor and $J_{\mu}$ an external current. The gauge-fixing term, with the parameter $\xi$, must be included in order to make the calculation of the propagator possible. In fact, as its known, without this term the wave operator would not be invertible. In this way, using the coefficients of corresponding wave operator,
$\alpha=-\frac{p^{2}}{\xi},\quad\beta=-p^{2}+\mu\,p^{4}\quad\text{and}\quad\gamma=i\kappa,$ (17)
@@
$\mathcal{O}^{-1}=-\frac{\xi}{p^{2}}P_{1}+\frac{1}{-p^{2}+\mu\,p^{4}+\kappa\sqrt{p^{2}}}P_{2}$
$\frac{1}{-p^{2}+\mu\,p^{4}-\kappa\sqrt{p^{2}}}P_{3}.$ (18)
Thus, at the basis $\{\omega,\theta,S\}$, we have
$\mathcal{O}^{-1}_{\mu\nu}=-\frac{p^{2}\left(1-\mu\,p^{2}\right)}{p^{2}\left[p^{2}\left(1-\mu\,p^{2}\right)^{2}-\kappa^{2}\right]}\,\theta_{\mu\nu}$
$\frac{i\kappa}{p^{2}\left[p^{2}\left(1-\mu\,p^{2}\right)^{2}-\kappa^{2}\right]}\,S_{\mu\nu}-\frac{\xi}{p^{2}}\,\omega_{\mu\nu}.$ (19)
@@ 5.3  The Maxwell-Deser-Jackiw Propagator
Another interesting model can be built up from the Maxwell system by adding the higher derivative Chern-Simons extension proposed by [5]
@@
At the basis $\{\omega,\theta,S\}$, with $\gamma=ip^{2}/m$, we have
@@ 6  General analysis of causality and unitarity
6.1  Analysis on the basis $\{P_{1},P_{2},P_{3}\}$
Let $J_{\mu}(p)$ be an external conserved current, $p^{\mu}J_{\mu}(p)=0$, the saturated amplitude is then
$\mathcal{A}(p)=J^{*\mu}(p)\,\mathcal{O}^{-1}_{\mu\nu}(p)\,J^{\nu}(p).$ (24)
Using the form (14), we restrict the analysis to the physical time-like region $p^{2}>0$, which corresponds to massive propagating modes. Space-like momenta ($p^{2}<0$) do not lead to physical on-shell excitations and therefore do not contribute to the unitarity analysis at tree level. In this regime we obtain
$\mathcal{A}(p)=-\frac{\|J^{(+)}\|^{2}}{\beta(p)-i\gamma(p)\sqrt{p^{2}}}-\frac{\|J^{(-)}\|^{2}}{\beta(p)+i\gamma(p)\sqrt{p^{2}}},$ (25)
where $J^{(+)}=P_{2}J$, $J^{(-)}=P_{3}J$ and the norms $\|J^{(\pm)}\|^{2}=-\,(J^{(\pm)})^{*\mu}\theta_{\mu\nu}J^{(\pm)\nu}\geq 0$ are taken in the physical transverse subspace. Thus, we will call the coefficient of each term in (25) as follows
@@ 6.1.1  Pole structure and general spectral analysis
Therefore, the poles associated with the vanishing of $D(p)$ are located at momentum $p^{\mu}=p_{*}^{\mu}$ satisfying
@@
For each fixed spatial momentum $\vec{p}$ these correspond to energies $p^{0}=E_{*}(\vec{p})$ given by the solutions of
@@
If $E_{*}(\vec{p})$ is real, we obtain a propagating mode with dispersion relation determined implicitly by $D(E_{*}(\vec{p}),\vec{p})=0$. For such modes, causality at the microscopic level requires that the group velocity
$v_{g}=\left|\frac{\partial E_{*}(\vec{p})}{\partial\vec{p}}\right|\leq 1,$ (32)
@@
$\frac{\partial D}{\partial p^{0}}\Big|_{p_{*}}\,dE_{*}(\vec{p})+\frac{\partial D}{\partial\vec{p}}\Big|_{p_{*}}\!\cdot d\vec{p}=0\quad\Rightarrow\quad\frac{\partial E_{*}(\vec{p})}{\partial\vec{p}}=-\,\frac{\partial D/\partial\vec{p}}{\partial D/\partial p^{0}}\Bigg|_{p_{*}}.$ (33)
@@
If some solutions $E_{*}(\vec{p})$ are complex, they occur in conjugate pairs by hermiticity of the underlying Lagrangian. Writing $E_{*}(\vec{p})=\Omega_{n}(\vec{p})-i\Gamma_{n}(\vec{p})$ with $\Gamma_{n}(\vec{p})>0$, the corresponding Fourier modes behave as $\exp[-i\Omega_{n}t]\,\exp[-\Gamma_{n}t]$ and can be interpreted this later on.
@@ 6.1.2  Residues and tree–level unitarity
$F_{\pm}(p^{0},\vec{p})\simeq\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}}\,\frac{1}{p^{0}-E_{*}(\vec{p})}.$ (36)
Therefore, the residues of the two contributions $\mathcal{A}_{\pm}$ to the saturated amplitude at $p^{0}=E_{*}(\vec{p})$ are
$\text{Res}\,\mathcal{A}_{\pm}\Big|_{p^{0}=E_{*}(\vec{p})}=-\|J^{(\pm)}\|^{2}\,\frac{\beta(p_{*})\pm i\gamma(p_{*})\sqrt{p_{*}^{2}}}{\left.\dfrac{\partial D}{\partial p^{0}}\right|_{p_{*}}},$ (37)
with $\|J^{(\pm)}\|^{2}\geq 0$. At this point it is convenient to adopt a more compact formulation. We reorganize the saturated amplitude so that the contributions associated with $D_{+}(p)$ and $D_{-}(p)$ appear explicitly,
$\mathcal{A}(p)=-\frac{D_{-}(p)\,\|J^{(+)}\|^{2}+D_{+}(p)\,\|J^{(-)}\|^{2}}{D(p)},$ (38)
where $D_{\pm}(p)=\beta(p)\mp i\gamma(p)\sqrt{p^{2}}$. Therefore, the residue of the saturated amplitude at $p^{0}=E_{*}(\vec{p})$ follows directly from the expansion of $D(p)$,
$\text{Res}\,\mathcal{A}\Big|_{p^{0}=E_{*}(\vec{p})}=-\frac{D_{-}(p_{*})\,\|J^{(+)}\|^{2}+D_{+}(p_{*})\,\|J^{(-)}\|^{2}}{\dfrac{\partial D}{\partial p^{0}}(p_{*})}.$ (39)
Since the pole condition implies either $D_{+}(p_{*})=0$ or $D_{-}(p_{*})=0$, the numerator simplifies accordingly. Indeed,
$D_{+}(p_{*})=0\quad$ $\Rightarrow$ $\quad D_{-}(p_{*})=2\beta(p_{*})=2i\gamma(p_{*})\sqrt{p_{*}^{2}},$
$D_{-}(p_{*})=0\quad$ $\Rightarrow$ $\quad D_{+}(p_{*})=2\beta(p_{*})=-2i\gamma(p_{*})\sqrt{p_{*}^{2}}.$
Thus, we obtain the following expressions for the residue
$\text{Res}\,\mathcal{A}\Big|_{p^{0}=E_{*}(\vec{p})}=-\frac{2\,\beta(p_{*})}{\partial_{0}{D}(p_{*})}\,\|J^{(\sigma)}\|^{2}=-\frac{\|J^{(\sigma)}\|^{2}}{\partial_{0}{D_{\sigma}}(p_{*})},\quad\sigma=\begin{cases}+,&D_{+}(p_{*})=0,\\ -,&D_{-}(p_{*})=0.\end{cases}$ (40)
Since $\|J^{(\pm)}\|^{2}\geq 0$, this means, according to the first expression, that the residue sign is determined by the sign of $\beta(p_{*})/\partial_{0}{D}(p_{*})$. For real poles, the saturated amplitude is real. In the models considered here, where $\gamma(p)$ is purely imaginary, each contribution is real separately, and no complex-conjugate pairing between the residues is required. Tree–level unitarity requires that the terms of (40) be non-negative.
6.1.3  Equivalent analysis in the basis $\{\omega,\theta,S\}$
@@
$\mathcal{A}(p)=J^{*}\!\cdot\mathcal{O}^{-1}\!J=\frac{\beta(p)}{D(p)}\,J^{*}\!\cdot J-\frac{\gamma(p)}{D(p)}\,J^{*}\!\cdot S\!J.$ (41)
@@
$J^{*}\!\cdot S\!J=(J^{(+)})^{*}\!\cdot SJ^{(+)}+(J^{(-)})^{*}\!\cdot SJ^{(-)}=$
$i\sqrt{p^{2}}\,\|J^{(+)}\|^{2}-i\sqrt{p^{2}}\,\|J^{(-)}\|^{2}.$ (44)
@@
$\mathcal{A}(p)=-\frac{1}{D(p)}\Bigl[\bigl(\beta(p)+i\gamma(p)\sqrt{p^{2}}\bigr)\,\|J^{(+)}\|^{2}\Bigr]$
@@
we recover exactly (25). Thus the decomposition and the analysis of poles, residues, unitarity and causality are entirely equivalent in the $\{P_{1},P_{2},P_{3}\}$ and $\{\omega,\theta,S\}$ basis. The projector formalism therefore provides a compact and geometrically transparent framework for studying higher–derivative planar gauge models at the level of their propagators.